OpenStudy (anonymous):

A bacteria population triples in number each day . if there are 2,551,500 bacteria on the 7th day , how many bacteria were present on the first day ?

OpenStudy (anonymous):

\[2,551,500=x\times 3^6\]

OpenStudy (anonymous):

\[x=\frac{2,551,500}{3^6}=3,500\]

OpenStudy (anonymous):

first day population is a a7 is population of 7th day a7=2551500 ratio is r=3 geometric progression sequence a7=a1 * r^n-1 n=7 2551500=a1 3^6 a1=2551500/729 =3500 population of bacteria on first day is 3500